OpenStudy (anonymous):

Identify the equations of ellipses whose major axis lengths are twice their minor axis lengths. 4x^2+25y^2+32x-250y+589=0 4x^2+y^2+16x+4y+4=0 16x^2+y^2-64x+8y+16=0 2x^2+8y^2-12x+16y-174=0 3x^2+12y^2+18x-24y-69=0 x^2+9y^2-2x+18y-71=0

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

sorry - ive forgotten this stuff

OpenStudy (anonymous):

ok:(

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@perl

OpenStudy (perl):

you want to find equations that have the two possible forms $$ \Large{ \frac{(x-h)^2}{(2a)^2} + \frac{(y-k)^2}{a^2}=1 \\~\\ \frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{(2a)^2}=1 } $$

OpenStudy (anonymous):

how? which is h and which is k?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

h depends on the particular equation, you can complete the square

OpenStudy (perl):

the long way is to complete the square for each equation

OpenStudy (perl):

to save time you can plug in x = 0

OpenStudy (perl):

here is the first graph https://www.desmos.com/calculator/kedo1nsysm

OpenStudy (perl):

if you hover your mouse you see that the major axis is length 1 - (-9) = 10 the minor axis is 7 - 3 = 4. and 10 is not twice 4. so that doesnt work

OpenStudy (perl):

hover your mouse over the vertices

OpenStudy (perl):

These three functions satisfy youre requirement: https://www.desmos.com/calculator/yhiapd89td

OpenStudy (perl):

actually looking at the standard form, we see

OpenStudy (perl):

$$ \Large{ \frac{(x-h)^2}{(2a)^2} + \frac{(y-k)^2}{a^2}=1 \\~\\\frac{(x-h)^2}{4a^2} + \frac{(y-k)^2}{a^2}=1 \\~\\ 4a^2\left( \frac{(x-h)^2}{4a^2} + \frac{(y-k)^2}{a^2}\right)=4a^2 \cdot 1 \\~\\ 4a^2\cdot \frac{(x-h)^2}{4a^2} + 4a^2 \cdot \frac{(y-k)^2}{a^2})=4a^2 \cdot 1 \\~\\ {(x-h)^2} + 4\cdot(y-k)^2=4a^2 \\~\\ {(x-h)^2} + 4\cdot(y-k)^2 -4a^2=0 }$$ similarly we can show $$\Large{ \frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{(2a)^2}=1 \\~\\ \iff \ \\~\\ 4(x-h)^2 + (y-k)^2 - 4a^2=0 } $$